OpenStudy (anonymous):

the coordinates are (1,3) , (2,5) , (3,6) , (4,9) , (5,10.5) , (6, 13) , (7,15) , (8,16) , and (9, 19) What is the equation of a line of best fit in slope-intercept form.

563blackghost (563blackghost):

Find the slope using the slope formula....\(\LARGE{\frac{ y2-y1 }{ x2-x1 }}\) Get two points and place them where they belong....so we would get the first two points... \(\LARGE{\frac{5-3}{2-1}}\) Once you get your answer that would be the slope in the slope-intercept equation.... \(\LARGE\color{green}{y=mx+b}\) \(\LARGE{~~~~~~~~|~~~~~~~~~|}\) \(\Large\color{purple}{~~~~~~slope~~y-intercept}\) To double check if your slope is correct you would get to other points and use the slope formula once again if they are alike then that is the offical slope... ~~~~~~~~~~~~~~~~~~~~~ To find the y-intercept you would plot your points and then draw a line between the points....when the line reaches the y-axis that would be the y-intercept....you will know what the y-intercept when x is 0....